A government agency was charged by the legislature with estimating the length of time it takes citizens to fill out various forms. Two hundred randomly selected adults were timed as they filled out a particular form. The times required had mean 12.8

 minutes with standard deviation1.7 minutes. Construct a 90%

 confidence interval for the mean time taken for all adults to fill out this form.


1
Expert's answer
2022-04-22T12:51:14-0400

"90\%CI=(12.8-1.652\frac{1.7}{\sqrt{200}},12.8+1.652\frac{1.7}{\sqrt{200}})=(12.601, 12.999)."